OpenStudy (anonymous):

How would I simplify 32/5? (its under a radical sign)

OpenStudy (hoblos):

\[\sqrt{32/5} = \sqrt{(4^{2}\times2)/5} = 4\sqrt{2/5}\]

OpenStudy (radar):

\[\sqrt{32/5}=\sqrt{32}\div \sqrt{5}=\sqrt{16\times2}\div \sqrt{5}=4\sqrt{2}\div \sqrt{5}\] \[4\sqrt{2}\over \sqrt{5}\]

OpenStudy (radar):

multiplying numerator and denominator by\[\sqrt{5}\] gives youy\[4\sqrt{10}\over5\]or \[0.8\sqrt{10}\]